OpenStudy (anonymous):

What is the solution set of |x – 2| + 4 = 3?

jimthompson5910 (jim_thompson5910):

|x – 2| + 4 = 3 |x – 2| + 4 - 4= 3 - 4 |x – 2| + 0 = -1 |x – 2| = -1

jimthompson5910 (jim_thompson5910):

That last equation, is it possible?

OpenStudy (anonymous):

What is the solution set of |x| < 15?

jimthompson5910 (jim_thompson5910):

Hint: If |x| < k, for some positive number k then -k < x < k

OpenStudy (anonymous):

x < 15 is my solution

jimthompson5910 (jim_thompson5910):

close

OpenStudy (anonymous):

really that is not the anwser

jimthompson5910 (jim_thompson5910):

no

OpenStudy (anonymous):

x > –15 and x < 15

jimthompson5910 (jim_thompson5910):

good, that's the same as saying -15 < x < 15

OpenStudy (anonymous):

ohh ty

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

An absolute value equation with two solutions An absolute value equation with one solution An absolute value equation with no solutions